I need help in graphing linear equations with fractions. For ex. y=1/2x-4. I dont understand how to find the values of x & y and how the problem is solved.
Found 2 solutions by rfer, richwmiller:
Answer by rfer(16322)

First of all the equation you posted is in the slope intercept form.
It has all the information you need to graph it.
Remember that -4 is the y intercept shown below.
y=mx+b
1/2 is the rise over run.
Start with x=0
y=1/2(0)-4
y=-4
(0,-4)
Now from there go right two and up one for next point, and so on.
Good luck.

Answer by richwmiller(17219)

To find the values of x and y
you pick one and find the other based on the equation
The easiest are when x=0 and when y=0
when x=0 y=-4
when y=0 x=2
you only need two points for a,linear equation because it forms a line .
You don't actually solve this equation. There are two unknowns .
you can solve the problem of graphing though.